OpenStudy (anonymous):

Kayla wants to use the Babylonian method to estimate ^43 to the nearest hundredth. Her initial estimate is 6.9. What is her estimate after she correctly completes the Babylonian method once? 1. 6.44 2. 6.45 3. 6.56 4. 6.57

OpenStudy (anonymous):

sorry dont know. hope u get it !

OpenStudy (anonymous):

that was a great answer and it deserves a medal

OpenStudy (anonymous):

@RED LMAOO Ook i ghess i meen this math is krazy nowadays!

OpenStudy (anonymous):

lol

OpenStudy (istim):

What's the Babylonian Method?

OpenStudy (istim):

Seriously? Don't gimme a medal until I actually do something.

OpenStudy (istim):

You're de-valuing the OS medals! There's gunna be an OS economical collapse!

OpenStudy (anonymous):

i didnt give anyone, any medals

OpenStudy (istim):

So, what's the Babylonian Method?

OpenStudy (anonymous):

its a method for computing square roots

OpenStudy (accessdenied):

The Babylonian method is a way of approximating square roots of numbers \[ x_{n+1} = \frac{1}{2}(x_{n} +\frac{S}{x_{n}}) \] Where S is the number under the square root and x_n is our first guess. \\ At least, this is what wikipedia tells me. I'm not familiar with this method myself. :P

OpenStudy (anonymous):

same here

OpenStudy (istim):

From what I see, we could possibly just sub in the values...

OpenStudy (accessdenied):

x_,n+1 ~ 6.5659... nearest hundreth being 6.57

OpenStudy (istim):

IS that what you need, Zalia?

OpenStudy (istim):

Imma go then. Good luck, and BATMAN AWAY

OpenStudy (anonymous):

lol k bye tim :)

OpenStudy (anonymous):

thnk u ^_^